If sin x = 0.32, Which of the following is sin(x +8π)?

A. 0.32
B. 0.64
C. 16.4
D. 164
1
Expert's answer
2013-06-06T10:16:16-0400
Sine function is periodic with period 2pi.
Due to it, sin(x+2pi*k)=sin x for any integer k.
In our case k=4: sin(x+8pi)=sin(x+4*2pi)=sin x
Thus, correct answer is A.